the future value of an investment of $10,000 at 0.4% per month compounded monthly after 17 years
Answer by checkley77(12569)

P=(1+r/n)^t*n Where P=principle, r=rate as a decimal, n= thr number if interest periods per year & t=the number of years.
10,000(1+.004/12)^17*12
10,000(1+.003333)^204
10,000(1.003333)^204
10,000*1.971513=$19,715.13 ans.